Assume that net income using the weighted-average cost flow assumption is $232,000. Calculate net income under FIFO and LIFOThe following data are available for Sellco for the fiscal year ended on January 31, 2017: Sales Beginning inventory Purchases, in chronological order 3,200 units 1,000 units 8 1,200 units10 1,600 units12 800 units 16 Required: a. Calculate cost of goods sold and ending inventory under the cost flow assumptions FIFO, LIFO and weighted average (using a periodic inventory system): (Round unit cost to 2 decimal places.) Cost Of Goods Sold Ending l FIFO LIFO Weighted averageS 32,000S 40,000 36.174| $ 20,000 12,000 15,826 b. Assume that net income using the weighted-average cost flow assumption is $232,000. Calculate net income under FIFO and LIFO. (Round unit cost to 2 decimal places.) Net Income FIFO LIFO
Answer #1

b. Calculation of net income under FIFO and LIFO method.

Since the net income under weighted average inventory method is given, the net income under FIFO and LIFO method can be found out by adjusting for the difference in cost of goods arising in each method.

As correctly calculated by the student above in a., the cost of goods sold under FIFO method is $4174 (36174-32000) less than the cost of goods sold under weighted average inventory method. To arrive at net income under FIFO method, we will have to add this difference to the net income under weighted average method.

Net income under FIFO method= 232000+4174= $236174

Similarly the cost of goods sold under LIFO method is $3826 (40000-36174) greater than the cost of goods sold under weighted average method. This has to be reduced from the net income under weighted average method to arrive at the net income under LIFO method.

Net income under LIFO method=232000- 3826= $228174